From: NASA HQ
Posted: Saturday, September 5, 2009
All ISS systems continue to function nominally, except those noted previously or below. Flight Day 9 of STS-128/17A. Saturday --- a very busy day for the crews.
Today’s sleep/wake cycle for the crew: Sleep – 3:30am; wake – 12:00pm; sleep tomorrow – 3:00am.
Tim Kopra & Nicole Stott completed another “liquid” session part of the biomed experiment INTEGRATED IMMUNE, collecting their liquid saliva samples right after wake-up. [Along with NUTRITION (Nutritional Status Assessment), INTEGRATED IMMUNE (Validating Procedures for Monitoring Crew member Immune Function) samples & analyzes participant’s blood, urine, and saliva before, during and after flight for changes related to functions like bone metabolism, oxidative damage and immune function to develop and validate an immune monitoring strategy consistent with operational flight requirements and constraints. The strategy uses both long and short duration crewmembers as study subjects. The saliva is collected in two forms, dry and liquid. The dry samples are collected at intervals during the collection day using a specialized book that contains filter paper. The liquid saliva collections require that the crewmember soak a piece of cotton inside their mouth and place it in a salivette bag; there are four of the liquid collections during docked operations. The on-orbit blood samples are collected right before undocking and returned on the Shuttle so that analysis can occur with 48 hours of the sampling. This allows assays that quantify the function of different types of white blood cells and other active components of the immune system. Samples are secured in the MELFI (Minus-Eighty Laboratory Freezer for ISS). Also included are entries in a fluid/medications intact log, and a stress-test questionnaire to be filled out by the subject at begin and end. Urine is collected during a 24-hour period, conventionally divided into two twelve-hour phases: morning-evening and evening-morning.]
Starting at ~1:45pm EDT, FE-1 Barratt & FE-4 Thirsk haven taken on the extensive 3.5 hr-task of removing & replacing the powered bolt 4-1 of the ACBM (Active Common Berthing Module) at the Node-2 nadir vestibule. As part of the job, the associated PCBM (Passive CBM) nut of the MPLM (Multi-Purpose Logistics Module) is also to be replaced. Later tonight, Frank will re-install the CPA-4 (Controller Panel Assembly 4) in the vestibule. [Without the R&R, the MPLM could not be unberthed.]
As the second important IFM (Inflight Maintenance) today, Barratt will join with FE-5 De Winne at ~9:00pm to continue troubleshooting of the OGA (Oxygen Generator Assembly System), performing the 3 hr-job of removing & replacing the Water ORU inlet screen. [This task involves removing one QD (Quick Disconnect), releasing the Manifold from the ORU, and cleaning all components prior to reassembly. Spare small O-ring seals were not found, but two small O-rings were recovered during pump filter R&R. One was likely exposed to catalyst and will not be re-used. If a spare is needed, the other will be evaluated to determine if it’s a viable spare.]
Mike is also timelined to relocate a PBA (Portable Breathing Apparatus, #1028) from the JLP (JEM Logistics Pressurized Segment), where it has become inaccessible due to stowage, to the Kibo JPM (JEM Pressurized Module).
FE-3 Romanenko conducted another spectrometer calibration session for the DZZ-12 RUSALKA (“Mermaid”) science payload, using solar observation from SM (Service Module) window #1. [RUSALKA ops involve calibration and tests of research equipment relating to the Sun and the Earth's limb at sunset (atmosphere lighted). To be tested are the procedure for remote determination of Methane (CH4) & Carbon Dioxide (CO2) content in the atmosphere (in the First Phase), measurement of CH4 & CO2 content in the atmosphere and reception of data on NI2 and NI4 content over the territories subjected to natural and technogenic effects, reception of sufficient data on seasonal dependencies of tropospheric parameters being studied (in the Second Phase). Equipment used: Rusalka monoblock, Nikon D2X digital photo camera(s); AF VR Nikkor ED 80-400f/4.5-5.6D lens with ultraviolet filter, bracket for attachment to the window, and a Rusalka-Accessories set.]
Continuing the current round of monthly preventive maintenance of RS (Russian Segment) ventilation systems, Romanenko is timelined to clean the four “Group B” fan screens (VT1, VTK1, VT2 & VVTK2) in the SM.
Similarly, CDR Padalka will do the periodic ventilation maintenance in the SM on the Group A fans and change out the cartridges of the four dust filters (PF1-4).
The CDR is also to “modernize” the Russian RSK2 laptop, installing a new shell (A31p laptop #1152 from US spares) to replace the failed shell and affixing Russian letter decals to laptop keyboard. Gennady will also move a new shell (31p #1157 from US spares) for the RS1 laptop to a stowage location.
Barratt, Kopra & Thirsk had more time set aside for more MPLM cargo transfers. Status of completion as of end-FD8:
Middeck – 98% (to ISS), 77% (from ISS)
MPLM – 91% (to ISS), 66% (from ISS)
Overall completed: 66%, on schedule.
FE-2 Kopra & FE-2-20 Stott are both scheduled for the periodic O-OHA (On-Orbit Hearing Assessment) test, a 30-min NASA environmental health systems examination to assess the efficacy of acoustic countermeasures, using a special software application on the MEC (Medical Equipment Computer) laptop. [The O-OHA audiography test involves minimum audibility measurements for each ear over a wide range of frequencies (0.25-10 kHz) and sound pressure levels, with the crewmembers using individual-specific Prophonics earphones, new Bose ANC headsets (delivered on 30P) and the SLM (sound level meter). To conduct the testing, the experimenter is supported by special EarQ software on the MEC, featuring an up/down-arrow-operated slider for each test frequency that the crewmember moves to the lowest sound pressure level at which the tone can still be heard. The baseline test is required not later than about Flight Day 14 for each new Expedition and is then generally performed every 45 days thereafter. Note: There have been temporary hearing deficits documented on some U.S. and Russian crewmembers, all of which recovered to pre-mission levels.]
Tim also has another hour set aside for regular crew departure preparations, working on the standard end-of-increment cleanup preparatory to his return to Earth on STS-128/17A. [It is usual for crewmembers to be granted reduced workdays for making their departure preparations, as their return date approaches.] Stott is to retrieve & stow the four passive FMK (Formaldehyde Monitoring Kit) sampling assemblies deployed by Mike Barratt 9/3 in the Lab (at P3, below CEVIS) and SM (at the most forward handrail, on panel 307), to catch any atmospheric formaldehyde on a collector substrate for subsequent analysis on the ground. [Two monitors each are usually attached side by side, preferably in an orientation with their faces perpendicular to the direction of air flow.] In addition, Nicole will conduct her first offloading of the WPA (Water Processor Assembly) from WRS (Water Recovery System) Rack 1 into a CWC-I (Collapsible Water Container-Iodine) with the common H2O transfer hose (which takes about 23 min) from the PWD (Potable Water Dispenser) Auxiliary Port, then flush the system.
Thirsk & De Winne are both on schedule for another session with the MedOps experiment WinSCAT (Spaceflight Cognitive Assessment Tool for Windows), Bob’s 5th and Frank’s 4th, by logging in on the MEC laptop and performing the psychological evaluation exercise on the PC-based WinSCAT application. [WinSCAT is a monthly time-constrained questionnaire test of cognitive abilities, routinely performed by astronauts aboard the ISS every 30 days before or after the PHS (periodic health status) test or on special CDR's, crewmembers or flight surgeons request. The test uses cognitive subtests that measure sustained concentration, verbal working memory, attention, short-term memory, spatial processing, and math skills. The five cognitive subtests are Coding Memory - Learning, Continuous Processing Task (CPT), Match to Sample, Mathematics, and Coding Delayed Recall. These WinSCAT subtests are the same as those used during NASA’s long-duration bed rest studies.]
Mike, Timothy, Nicole & Bob have time set aside for filling out their regular weekly FFQ (Food Frequency Questionnaire) on the MEC (Medical Equipment Computer). [On the FFQs, NASA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. The FFQ is performed once a week to estimate nutrient intake from the previous week and to give recommendations to ground specialists that help maintain optimal crew health. Weekly estimation has been verified to be reliable enough that nutrients do not need to be tracked daily.]
The CDR will do the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]
The FE-3 is scheduled for the daily IMS (Inventory Management System) maintenance, updating/editing its standard “delta file” including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).
Afterwards, Roman conducts the periodic checkout & performance verification of IP-1 airflow sensors in the various RS hatchways.
FE-5 De Winne will take documentary photography of the three JAXA MDS (Microbial Detection Sheet) devices used by him yesterday to take surface samples in the Kibo JPM (Saibo Rack wall, Node-2/JEM hatch handrail, JPM OA1 diffuser), then stow the samples in the MELFI-1 (Minus Eighty Laboratory Freezer for ISS, Dewar 4, Tray A, Section 4).
Frank also has an hour reserved to continue troubleshooting the ROBoT simulator program, doing performance test runs of the application with the new uplinked software patch. [After the troubleshooting, Frank was to clean up any old patches and save the uplinked patch to USB devices.]
The ISS crew are performing their regular daily 2.5-hr. physical workout program on the CEVIS cycle ergometer (FE-1, FE-2-20, FE-4, FE-5), TVIS treadmill with vibration isolation (CDR, FE-2, FE-3), ARED advanced resistive exercise device (FE-1, FE-2, FE-4, FE-5), and VELO cycle ergometer with bungee cord load trainer (CDR).
Afterwards, Frank De Winne transfers the exercise data files to the MEC for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on ARED, followed by their erasure on the HRM storage medium (done six times a week).
When the ground has powered down the external DDCU (DC-to-DC Converter Unit) S01A tonight during the EVA-3 for the RPCM (Remote Power Controller Module) R&R, the FE-4 will use this opportunity to safe an ungrounded floating connector in the Lab (loc. LAB1P3) by covering it with a CTB (Cargo Transfer Bag) divider and gray tape.
The usual pre-EVA activities like EMU batteries recharge, D2Xs camera turn-around, amateur/ham radio power down, protective shutters closing for Lab & JPM windows, etc. will be handled by Barratt, Stott and De Winne.
PMA-3 Heater Cable Installation: After the EVA-2 crew discovered that the PMA-3 (Pressurized Mating Adapter 3) is rotated (“clocked”) 90 degrees off (“tail nadir”) from the expected orientation (”tail aft”), further evaluation by ground teams confirmed that the on-board jumper harness is not long enough in this orientation. Therefore, this task will not be performed during EVA-3. It was also confirmed that the actual PMA-3 orientation is the same as in 2007 (Expedition 15) before it was relocated to Node-1 nadir, and that this is the correct orientation expected for HTV rendezvous/berthing. There are currently no plans to rectify the situation which requires that entrances to PMA-3 need to be restricted to certain (solar-heated) Beta angle periods to prevent condensation within the module. The root cause for this discrepancy is under investigation.
Buzz Lightyear Video: During the EVA ingress tonight, Tim Kopra will be taking high-definition video of the Buzz Lightyear doll for an "EVA" Episode. [This video, which should capture footage of crew ingressing from an EVA, is part of the ISS National Lab Office partnership with Disney/Pixar. It will be used in educational and outreach videos produced by Pixar, with a “Toy Story” movie spin, to stimulate young children's interest in NASA and STEM (Science, Technology, Engineering, and Math) education.]
After Danny Olivas (EV1) & Christer Fuglesang (EV2) spent last night on “campout” in the U.S. Airlock (A/L) on 10.2 psi atmospheric pressure, their EVA-3 activities today are scheduled as follows:
Post-ingress activities by Olivas, Fuglesang, Stott, Kopra & Hernandez will include the usual post-EVA tasks like photographing EMU gloves for inspection, recharging EMUs with water, downloading & downlinking D2XS EVA & glove photographs, recharging REBA batteries, etc.
Sleeptime for the crews will begin tomorrow morning at 3:00am, to last until 11:30pm.
3-D SPACE: “Tim, thanks for the performance of your final session on 09/02 (GMT245)! Thanks also for configuring the Digital Tablet pens for future use.”
AgCam (Agricultural Camera): No report.
ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.
ALTEA DOSI (NASA/ASI): Standing by.
BCAT-4/5 (Binary Colloidal Alloy Test 4/5): No report.
BIOLAB (ESA): No report. Biological Rhythms (JAXA, BIORHYTHMS): No report.
BISE (CSA, Bodies in the Space Environment): No report.
BISPHOSPHONATES: “Bob, thanks for completing your pill ingestion. Your next session is scheduled for 9/7”. CARD (Long Term Microgravity: Model for Investigating Mechanisms of Heart Disease, ESA): No report.
CB (JAXA Clean Bench): No report.
CBEF (JAXA Cell Biology Experiment Facility)/SPACE SEED: CBEF Fan Sensor removal was completed on 8/28. After few CBEF Fan repairs and the checkout, we will start the Space Seed experiment from 9/10. Space Seed KFT’s were inserted into MELFI prior to the experiment star. Due to an error in stowage note, the actuator KFT’s were inserted into MELFI instead of the Space Seed KFT’s. However, this was corrected the next day and the correct KFT’s have been successfully inserted into MELFI.
CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): “Tim, thanks for completing your last on-orbit CCISS session. We look forward to seeing you on the ground.”
DOSIS (ESA): Experiment is progressing nominally with active and passive dosimeters measurements. Data downlink was performed on 8/31.
EarthKAM (Earth Knowledge Acquired by Middle School Students): Planned.
EDR (European Drawer Rack, ESA): The rack is continuously active in support of the Protein Crystallization Diagnostic Facility (PCDF) experiment. EDR is providing power/data and temperature control (via cooling loop) to PCDF.
ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.
ENose (Electronic Nose): No report.
EPM (European Physiology Module): MEEMM HD was packed on 8/26 in order to be returned on 17A (STS-128). EPO (Educational Payload Operations, NASA): No report.
EPO 3-min Video (JAXA): No report. EPO J-Astro Report (JAXA): Ongoing. EPO Space Clothes (JAXA): Complete.
EPO Hiten (Dance, JAXA): No report.
EPO Moon Score (JAXA): No report.
EPO Try Zero-G (JAXA): “No report.
EPO Kibo Kids Tour (JAXA): Complete. EPO Spiral Top (JAXA): No report.
ETD (Eye Tracking Device): Completed.
EuTEF (European Technology Exposure Facility): On 9/1 during 17A EVA-1, the EuTEF platform and its instruments have been de-installed from the CEPF (Columbus External Platform Facility) and successfully transferred to the Shuttle cargo bay for return. For some instruments, science acquisition was performed until a couple of hours prior to de-installation.
FACET (JAXA): No report.
FSL (Fluid Science Laboratory): No report.
GEOFLOW: No report.
HDTV System (JAXA): To be launched by HTV1.
Holter ECG (JAXA): No report.
HQPC (JAXA): To be launched by 34P.
ICE CRYSTAL (JAXA): Complete.
ICV (Integrated Cardiovascular): “Nicole, the ICV team is looking forward to your first echo session next week. Bob and Frank have done this three times already with wonderful results so you are in excellent hands. Welcome aboard!” IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.
INTEGRATED IMMUNE: No report.
InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): No report.
IRIS (Image Reversal in Space): No report.
LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): No report.
MEIS (Marangoni Experiment for ISS) in JAXA FPEF (Fluid Physics Experiment Facility): “16 runs completed. MAXI (Monitor of All-sky X-ray Image, JAXA): On 8/25, all MAXI data transmission though Ethernet had stopped. Trouble shooting was completed. After power cycles, Ethernet and other communication has returned.
MDCA/Flex: See under CIR.
Microbe-1 (JAXA): Microbe Swab sampling on the 3 diffusers and the tube installation into MELFI has been completed.
Micro-G Clay (JAXA EPO): Complete. MMA (JAXA/Microgravity Measurement Apparatus):MMA data downlink performed on 7/1. FSL Optical Target #4 check-out was completed by 7/1.
NEUROSPAT (ESA/Study of Spatial Cognition, Novelty Processing and Sensorimotor Integration): No report. NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete. NUTRITION w/REPOSITORY: Ongoing. PADLES (JAXA, Area PADLES 3; Passive Area Dosimeter for Lifescience Experiment in Space): Area Dosimeters (PADLES 2) were removed for return on 17A.
PCDF-PU (Protein Crystallization Diagnostic Facility - Process Unit): No report.
PCG (JAXA, Protein Crystal Growth): Temperature monitoring by the ground team is now in progress. Experiment continues to be in good condition. PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete. PMDIS (Perceptual Motor Deficits in Space): Complete.
SEDA-AP (Space Environment Data Acquisition Equipment-Attached Payload, JAXA): All monitor functions have been verified. Temperature monitoring by ground operation is now in progress.
SHERE (Shear History Extensional Rheology Experiment): Complete. SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): The next monthly Sleep Actiwatch download will be on the task list from 9/12 to 9/18 and hard scheduled on 9/20. The Actiwatches will stop recording data on 9/20.
SMILES (JAXA): Exposed Payload, to be launched by HTV1. SOLAR (Solar Monitoring Observatory): The current Sun observation window ended on 0/25. Next Sun observation window is predicted to start on 9/15.
SOLO (Sodium Loading in Microgravity): No report. SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): No report.
SPICE (Smoke Point In Co-flow Experiment): No report.
SWAB (Characterization of Microorganisms & Allergens in Spacecraft): No report.
TRAC (Test of Reaction & Adaptation Capabilities): Planned.
VLE (Video Lessons ESA): VLE-1 completed.
WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete/Planned (2J/A Stage). No report.
CEO (Crew Earth Observations): “Through 9/1 the ground has received a total of 26,005 frames of ISS CEO imagery for review and cataloging. “We are continuing to work through a sizeable backlog of your plentiful imagery. We are pleased to confirm your acquisition of imagery for the following CEO target requests: Heidelberg, Germany; Haruj Megafans, Libya – under review; Black Point Lava Flow, Arizona – excellent views – requirements complete; Hurricane Bill, Atlantic Ocean; Lake Poopo, Bolivia – excellent detail, incomplete; Toshka Lakes, Egypt – under review; All Saints Bay, Salvador, Brazil – you found parts of it despite the clouds; and Sao Paulo-Rio de Janeiro Aerosol – under review. Nice work! Your recent image of the Heiltskuk Ice Field in the Coast Mountains of British Columbia, Canada was published on NASA/GSFC’s Earth Observatory website this past weekend. This image is one of the best we have ever seen of this long standing CEO target area and provides excellent baseline imagery for changes in the glaciers of this spectacular ice field. Great Shot!”
No CEO photo targets uplinked for today. CEO photography can be studied at this “Gateway” website: http://eol.jsc.nasa.gov (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).
STS-128/Discovery -17A Crew & Mission Timeline: v CDR: Fredrick W. “Rick” Sturckow v PLT: Kevin A. Ford v MSs: Patrick G. Forrester; Jose M. Hernandez; John D. “Danny” Olivas; Christer Fuglesang v ISS FE-2s: Nicole Passonno Stott (UP); Timothy L. Kopra (DOWN) . FD09 (9/05) -- EVA3; deploy S3 PAS, R&R RGA 2, connect PMA-3 heater cable, R&R S0 RPCM, route Node-3 cabling . FD10 (9/06) -- Crew off duty; final cargo transfers; initiate O2 transfer . FD11 (9/07) -- Egress & uninstall MPLM; berth MPLM in PLB, terminate O2 transfer, close hatches . FD12 (9/08) -- Undock from ISS (~3:27pm EDT); perform flyaround; dump H2O; late inspection . FD13 (9/09) -- Cabin stow; Orbiter FCS checkout; RCS hot fire; H2O dump . FD14 (9/10) -- Nominal deorbit (6:05pm); landing (~7:08pm KSC).
ISS Crew Sleep Shifting: To synchronize the ISS crew’s timeline with STS-128/17A arrival and docked activities, the station’s wake/sleep cycle is undergoing a number of shifts which started on 8/29 and ends on 9/9. For the next few days, the schedule is as follows: